# A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6

### v1

@@ 4.1 Polarisation fraction and morphology of magnetic fields
Owing to the wavelength coverage and the spatial resolution provided by ALMA, we can observe similar rest-frame FIR and reach comparable spatial resolutions as the dust polarisation observations in the nearby Universe, especially compared with what SOFIA/HAWC+ has done in the SALSA Legacy Program (Lopez-Rodriguez et al. 2022b). In Fig. 4, we compare the measured dust polarisation fraction in SPT0346-52 with the measurements from the SALSA Legacy Program. The rest-frame wavelength probed in SPT034652 is at 131 m, averaged over a physical resolution about 700 pc, comparable to the spatial resolution of SALSA Legacy Program at 154 m (250 pc-1.5 kpc, Lopez-Rodriguez et al. 2022b). A typical mean resolved polarisation fraction of per cent has been reported among all the galaxies in the SALSA survey, dropping to a lower fraction of per cent if considering only starburst galaxies (SFR). We also include the dust polarisation measurements of 9io9 (Geach et al. 2023). The similar polarisation fraction between SPT034652 and the local starbursts is surprising, even though there are two orders of magnitude differences in SFR and more than 12 Gyr difference in cosmic time.
In SPT034652, the polarised dust emission is more patchy. We only observe polarised dust emission in a part of the galaxy, and interestingly, it does not overlap with the peak of total dust intensity (see Fig. 2). In Fig. 5, we compare the spatial distribution of polarisation with the kinematics from [C ii]. The region with the polarisation detection is also the region with the lowest velocity gradient and velocity dispersion. In the channel maps of [C ii] emission, the polarisation is mostly associated with a single velocity component near . This result is different from the configuration of polarised dust emission in 9io9, where the polarised dust is largely overlapped with the total dust emission and the inferred fields line are aligned with cold molecular gas disk (Geach et al. 2023). Such a difference suggests a possible different ordering process (see more discussion in section 4.2).
@@ 4.2 Physical origin of the ordered magnetic field
Various processes could lead to the formation of large-scale ordered magnetic fields, including the dynamo in galactic disks (Beck et al. 1994; Beck 2007), the stretching and compressing of existing fields by galaxy mergers, and galactic inflows/outflows (Chyży & Beck 2004; Drzazga et al. 2011; Basu et al. 2017; Jones et al. 2019b; Lopez-Rodriguez et al. 2022a; Lopez-Rodriguez 2023). In this section, we will discuss several possible mechanisms to explain the observed magnetic fields in SPT034652.
@@ 4.2.1 Disk or merger driven?
The formation of disks is the most discussed channel to develop galactic ordered magnetic fields (Beck et al. 1994; Pakmor et al. 2017, 2024). The presence of a galactic disk can provide continuous differential rotation to support the dynamo (Beck et al. 2019). Initially, dynamo was not favoured in the early Universe, as most of the early galaxies were thought to be irregular and turbulent. This picture may need to be revised, however, given the apparent prevalence of disks in the early Universe (Förster Schreiber et al. 2018; Rizzo et al. 2023; Kuhn et al. 2023, Lee et al. 2024 (in prep.)). If galaxies can develop disks efficiently during the first Gyr of the cosmic time, it is reasonable to expect large-scale ordered magnetic fields in the disks of early galaxies (Arshakian et al. 2009; Geach et al. 2023).
@@
where the is the scale height of the disk, is the angular velocity of the disk; and is the maximum velocity and velocity dispersion of the disk; is the maximum radius of the disk. Within a galactic disk, the dynamo is effective only when the dynamo number is larger than the critical value , which is close to 7 in Milky Way-like disk galaxies. The from our lens model is about 1.5 kpc (see also Fig. 3). Assuming a typical disk scale height of 300 pc (Geach et al. 2023), the is in the order of . However, considering the complex velocity field (merger+outflows), it is not straightforward to estimate the without some biased assumptions. In the most extreme case, assuming the velocity gradient is solely produced by a galactic disk, the maximum rotational velocity will be km s-1. We adopted the median velocity dispersion of 140 km s-1 from the region with the lowest velocity gradient, which is also the region with polarisation detection (see also Fig. 5). Then, the is about 2.5, which predicts , well below the (Ruzmaikin et al. 1988; Arshakian et al. 2009, see also the discussion in Geach et al. (2023)). Therefore, even if there is a disk in SPT034652, it will not be very effective in ordering the galactic magnetic fields. However, if the dynamic cold disk () could exist at such early cosmic time (e.g. Rizzo et al. 2020; Neeleman et al. 2020; Lelli et al. 2021; Rowland et al. 2024), the dynamo number could be large enough to allow efficient dynamo.
@@
Galaxy mergers can also compress and stretch magnetic fields, creating large-scale ordered magnetic fields. This process has been well demonstrated in the local major merger prototype – NGC4038/39 (the Antennae) (Chyży & Beck 2004; Basu et al. 2017). In the cores of the two merger galaxies of the Antennae system, the field lines are mostly dominated by turbulent fields because of the enhanced star formation. However, in the tidal tail, the magnetic fields are regular up to kpc, which is likely generated by stretching the initial disk fields in the pre-merger galaxies (Basu et al. 2017). A similar 9 kpc ordered magnetic field is also reported by means of magnetically aligned dust grains, which also supports a compressed and stretched magnetic field due to the merger activity (Lopez-Rodriguez et al. 2022a). The presence of ordered magnetic fields has also been confirmed in a sample of interacting galaxies observed with synchrotron polarisation (at 4.86 and 1.4 GHz) (Drzazga et al. 2011), which indicates tidal interactions can effectively magnetise the mergers’ surroundings and produce large-scale ordered field structures in the outskirts of the merged galaxies. Besides, the developed kpc-scale ordered magnetic fields have a strength several times larger than the magnetic fields in disks from non-interacting galaxies (Drzazga et al. 2011; Beck & Wielebinski 2013). Similar intense field amplification and creation of large-scale magnetic fields with galaxy mergers have also been noticed in magneto-hydrodynamic simulations, which is crucial for angular momentum transportation and subsequent morphological evolution of the merged galaxies (Whittingham et al. 2021, 2023).
@@
In SPT034652, the polarised dust emission is offset from the total dust emission, which is similar to the polarised dust distribution in Antennae galaxies (Lopez-Rodriguez et al. 2022a). In addition, the bimodal distribution of also favour the merger scenario. The position angles of the field lines depend only on the ratio of Stokes and , which have received identical amplification and averaging from the strong lensing. The presence of two groups of in the image plane indicates that the source plane should also have at least two magnetic field components, i.e., and in Fig. 2. For a merging system like the Antennae galaxies, it could easily develop multiple-component fields, i.e. the existing fields on each galaxy, the fields connecting both galaxies, and the newly stretched lines in the tidal tail. Alternatively, a disk could also show a bimodal distribution of , but it needs to be highly inclined towards the sightline (such as NGC7331 in Borlaff et al. 2023), and it cannot explain why the polarised dust mostly overlapped with a single velocity component. The compact size of SPT034652 favours a late-stage merger. It supports the possibility that galactic magnetic fields have already been established in the two merging galaxies in SPT034652, and the merger-triggered starburst makes the B-field more turbulent and leaves mainly the ordered magnetic fields in the outskirts. Therefore, the galaxy merger provides the most natural explanation for the formation of ordered magnetic fields in SPT034652.
@@ 4.2.2 Outflow driven?
A strong outflow is present in SPT034652 based on the high-resolution dust continuum and H2O() observations (Jones et al. 2019a). It has a velocity up to 500 km s-1 and a substantial outflow rate, 100-900 M⊙ yr-1, likely driven by the galactic wind powered by the compact starbursting core. The galactic wind can distribute amplified magnetic fields to a larger volume (Brandenburg et al. 1993; Arámburo-García et al. 2021), thus can also explain the ordered field in the outskirts of SPT034652. The best observational example can be found in the bipolar outflow of M82, where the geometry of the field lines are perfectly aligned with the outflow (Jones et al. 2019b; Lopez-Rodriguez et al. 2021). The strength of ordered magnetic fields is also strong in the outflow, which is about two orders of magnitude stronger than the ordered fields in the galactic disks (Lopez-Rodriguez et al. 2021).
@@ 4.2.3 AGN driven?
It is also worth noting the possible contribution from the AGN in SPT034652. Powerful jets from AGN play an important role in magnetising the intracluster medium (ICM, Dubois et al. 2009; Xu et al. 2009). Meanwhile, the large-scale shocks induced by the jet and the AGN-driven outflow could also produce large-scale shear to order the turbulent fields. In the less extreme radio-quiet AGNs, the ubiquitous presence of compact radio jets (e.g. Girdhar et al. 2022; Rivera et al. 2023) also indicates their possible contribution to the development of magnetic fields in the ISM. In SPT034652, however, the presence of AGN is still elusive. The spectrum energy distribution (SED) analysis suggests a negligible AGN contribution to the total energy output, which is supported by the non-detection of X-ray and faint radio emission (Ma et al. 2016). In nearby AGN host galaxies, only the radio-loud AGN show highly polarised FIR fraction (5–11 per cent) in the core region; the radio-quiet AGN are mostly unpolarised (Lopez-Rodriguez et al. 2023). Therefore, our results support that the possible embedded AGN in SPT034652 plays a negligible role in ordering the galactic magnetic fields. Future surveys targeting the AGN host galaxies are needed to test AGN’s importance in magnetising the ISM.
@@ 4.2.4 Caveats
In summary, a merger event provides the most likely explanation of the ordered magnetic fields in SPT034652, but we cannot fully rule out the possible contribution from outflows and a compact galactic disk. It is also possible that all these processes may work simultaneously to amplify and order the magnetic fields. Therefore, future high-resolution follow-ups will be the key to distinguish different scenarios, alleviate the beam averaging.
@@ 4.3 Formation timescales of galactic magnetic fields
The detection of the dust polarisation in SPT034652 indicates the development of kpc-scale ordered galactic magnetic fields in the Universe is less than 1 Gyr, which puts a strong constrain on the ordering timescale. Theoretically, the amplification via small-scale dynamos is efficient, capable of amplifying the seed magnetic fields to G turbulent magnetic within a timescale of 0.1 Gyr (Schleicher et al. 2010; Pakmor et al. 2014; Rieder & Teyssier 2016; Martin-Alvarez et al. 2022). However, the large-scale dynamos may need timescales of several Gyr to create kpc-scale ordered fields (e.g. Wang & Abel 2009; Kotarba et al. 2009; Arshakian et al. 2009; Rodrigues et al. 2019). Therefore, our results suggest that the ordering process from large-scale dynamos is more efficient in the early Universe than previously predicted. Meanwhile, the current two measurements of dust polarisation at , 9io9 and SPT0346-52, are both extreme starburst galaxies, which have order-of-magnitude higher SFR and gas fractions than the local SFGs. Such large SFR and high gas fraction normally indicate a highly turbulent ISM (Elmegreen & Scalo 2004; Bournaud et al. 2010), which may lead to the dominance of turbulent magnetic fields. If what we found in this new galaxy is a large-scale ordered B-field from a merger activity, then we would expect many more dusty SFGs in the early Universe to also have detectable polarised dust emission, which can be tested by future ALMA surveys.
@@
where is the volume density of the ISM, and is the velocity dispersion of the gas. In SPT0346-52, the observed [C ii] line dispersion at the region with polarisation detection is about 140 km s-1 (see also Fig. 5). We assume a preferred gas density of 50 cm-3 from the photodissociation regions (PDR) modelling (Litke et al. 2022). We estimate the equipartition turbulent magnetic field strength to be G. We emphasise that this is only a theoretical upper limit for the turbulent magnetic fields, with the assumption that the turbulence dominates the line width. If the line width is dominated by regular rotation or the velocity difference of the possible two merging galaxies, the equipartition field strength can be several times lower, which will bring the field strength closer to the local starburst galaxies, either measured with dust polarisation (Lopez-Rodriguez et al. 2021; Lopez-Rodriguez 2023) and radio synchrotron observations (Lacki & Beck 2013). It should be noted that the magnetic field strength in the cold phase of the interstellar medium is different from that of the warm and diffuse medium measured at radio wavelengths (Martin-Alvarez et al. 2023). The field strength in the cold and dense medium could be larger than that from the warm and diffuse medium, as the field strength correlates with the gas volume density. Nonetheless, our calculation supports that highly amplified magnetic fields could be developed in very young star-forming galaxies.
Various statistical evidence supports the claim that the early galaxies could have widely developed amplified magnetic fields. As we have mentioned before, the statistical excess of Faraday rotation in the quasar absorber and fast radio burst all support a wide presence of coherent magnetic fields in galaxies across (Bernet et al. 2008, 2013; Kronberg et al. 2008; Farnes et al. 2014, 2017; Mannings et al. 2023). Besides, in local star-forming galaxies, the total magnetic field strength correlates with the SFR surface density, i.e. with to be in the range of 0.14-0.3 (e.g. Beck 2015), together with the tight FIR-radio correlation (Condon 1992; Bell 2003), it betrays the close bond between the magnetic fields and star formation activities. In SPT034652, the SFR surface density is about , supporting a highly turbulent ISM with strongly amplified turbulent magnetic fields. Meanwhile, the FIR-radio correlation has been observed up to by stacking analysis, and it shows only mild variations with the redshift (e.g. Ivison et al. 2010; Magnelli et al. 2015; Delvecchio et al. 2021), suggesting that amplified magnetic fields may already be present in the star-forming galaxies up to . All the evidence supports the claim that small-scale dynamo efficiently amplifies the turbulent magnetic fields in the early star-forming galaxies. Fig. 4 also shows the magnetic field strength as a function of redshift and total star-formation rates. Due to the selection effects, our results are biased to the most extremely SFGs in the early Universe.

### v2

@@ 4.1 Polarisation fraction and morphology of magnetic fields
Owing to the wavelength coverage and the spatial resolution provided by ALMA, we can observe similar rest-frame FIR and reach comparable spatial resolutions as the dust polarisation observations in the nearby Universe, especially compared with what High-resolution Airborne Wideband Camera Plus (HAWC+) has done in the SALSA Legacy Program (Lopez-Rodriguez et al. 2022b). In Fig. 4, we compare the measured dust polarisation fraction in SPT0346-52 with the measurements from the SALSA Legacy Program. The rest-frame wavelength probed in SPT034652 is at 131 m, averaged over a physical resolution about 700 pc, comparable to the spatial resolution of SALSA Legacy Program at 154 m (250 pc-1.5 kpc, Lopez-Rodriguez et al. 2022b). A typical mean resolved polarisation fraction of per cent has been reported among all the galaxies in the SALSA survey, dropping to a lower fraction of per cent if considering only starburst galaxies (SFR). We also include the dust polarisation measurements of 9io9 (Geach et al. 2023). The similar polarisation fraction between SPT034652 and the local starbursts is surprising, even though there are two orders of magnitude differences in SFR and more than 12 Gyr difference in cosmic time.
In SPT034652, the polarised dust emission is more patchy. We only observe polarised dust emission in a part of the galaxy and, interestingly, it does not overlap with the peak of total dust intensity (see Fig. 2). In Fig. 5, we compare the spatial distribution of polarisation with the kinematics from [C ii]. The region with the polarisation detection is also the region with the lowest velocity gradient and velocity dispersion. In the channel maps of [C ii] emission, the polarisation is mostly associated with a single velocity component near . This result is different from the configuration of polarised dust emission in 9io9, where the polarised dust is largely overlapped with the total dust emission and the inferred fields line are aligned with cold molecular gas disk (Geach et al. 2023). Such a difference suggests a possible different ordering process (see more discussion in section 4.2).
@@ 4.2 Physical origin of the ordered magnetic field
Various processes could lead to the formation of large-scale ordered magnetic fields, including the dynamo in galactic disks (Beck et al. 1994; Beck 2007), the stretching and compressing of existing fields by galaxy mergers, and galactic inflows/outflows (Chyży & Beck 2004; Drzazga et al. 2011; Basu et al. 2017; Jones et al. 2019b; Lopez-Rodriguez et al. 2022a; Lopez-Rodriguez 2023; Ledos et al. 2024). In this section, we will discuss several possible mechanisms to explain the observed magnetic fields in SPT034652.
@@ 4.2.1 Disk or merger driven?
The formation of disks is the most discussed channel to develop galactic ordered magnetic fields (Beck et al. 1994; Pakmor et al. 2017, 2024). The presence of a galactic disk can provide continuous differential rotation to support the dynamo (Beck et al. 2019). Initially, dynamo was not favoured in the early Universe, as most of the early galaxies were thought to be irregular and turbulent. This picture may need to be revised, however, given the apparent prevalence of disks in the early Universe (Förster Schreiber et al. 2018; Rizzo et al. 2023; Kuhn et al. 2023, Lee et al. (in prep.)). If galaxies can develop disks efficiently during the first Gyr of the cosmic time, it is reasonable to expect large-scale ordered magnetic fields in the disks of early galaxies (Arshakian et al. 2009; Geach et al. 2023).
@@
where the is the scale height of the disk, is the angular velocity of the disk; and is the maximum velocity and velocity dispersion of the disk; is the maximum radius of the disk. Within a galactic disk, the dynamo is effective only when the dynamo number is larger than the critical value , which is close to 7 in Milky Way-like disk galaxies. The from our lens model is about 1.5 kpc (see also Fig. 3). Assuming a typical disk scale height of 300 pc (Geach et al. 2023), the is in the order of . Considering the complex velocity field (merger+outflows), it is not straightforward to estimate the without some biased assumptions. In the most extreme case, assuming the velocity gradient is solely produced by a galactic disk, the maximum rotational velocity will be km s-1. We adopted the median velocity dispersion of 140 km s-1 from the region with the lowest velocity gradient, which is also the region with polarisation detection (see also Fig. 5). Then, the is about 2.5, which predicts , well below the (Ruzmaikin et al. 1988; Arshakian et al. 2009, see also the discussion in Geach et al. (2023)). Therefore, even if there is a disk in SPT034652, it will not be very effective in ordering the galactic magnetic fields. However, if the dynamic cold disk () could exist at such early cosmic time (e.g. Rizzo et al. 2020; Neeleman et al. 2020; Lelli et al. 2021; Rowland et al. 2024), the dynamo number could be large enough to allow efficient dynamo.
@@
Galaxy mergers can compress and stretch magnetic fields, creating large-scale ordered magnetic fields. This process has been well demonstrated in the local major merger prototype – NGC4038/39 (the Antennae) (Chyży & Beck 2004; Basu et al. 2017). In the cores of the two merging galaxies of the Antennae system, the field lines are mostly dominated by turbulent fields. However, in the tidal tail, the magnetic fields are regular up to kpc, which is likely generated by stretching the initial disk fields in the pre-merger galaxies (Basu et al. 2017). A 9 kpc ordered magnetic field is also reported by means of magnetically aligned dust grains, which also supports a compressed and stretched magnetic field due to the merger activity (Lopez-Rodriguez et al. 2022a). The presence of ordered magnetic fields has also been confirmed in a sample of interacting galaxies observed with synchrotron polarisation (at 4.86 and 1.4 GHz) (Drzazga et al. 2011), which all support tidal interactions can effectively magnetise the mergers’ surroundings and produce large-scale ordered field structures up to tens of kpc in the ISM of the merged galaxies. Besides, the developed kpc-scale ordered magnetic fields have a strength several times larger than the magnetic fields in disks from non-interacting galaxies (Drzazga et al. 2011; Beck & Wielebinski 2013). Similar intense field amplification and creation of large-scale magnetic fields with galaxy mergers have also been noticed in magneto-hydrodynamic simulations, which is crucial for angular momentum transportation and subsequent morphological evolution of the merged galaxies (Whittingham et al. 2021, 2023).
@@
In SPT034652, the polarised dust emission is offset from the total dust emission, similar to the polarised dust distribution in Antennae galaxies (Lopez-Rodriguez et al. 2022a), which is likely due to the disruption of ordered fields by the merger-trigger starbursts. In addition, the bimodal distribution of also favour the merger scenario. The position angles of the field lines depend only on the ratio of Stokes and , which have received identical amplification and averaging from the strong lensing. The presence of two groups of in the image plane indicates that the source plane should also have at least two magnetic field components, i.e., and in Fig. 2. For a merging system like the Antennae galaxies, several configurations of magnetic field can be formed, i.e. the existing fields on each galaxy, the fields connecting both galaxies, and the newly stretched lines in the tidal tail. Alternatively, a disk could also show a bimodal distribution of , but it needs to be highly inclined towards the sightline (such as NGC7331 in Borlaff et al. 2023), and it cannot explain why the polarised dust mostly overlapped with a single velocity component. Besides, the dynamo predicts regular magnetic fields with axisymmetric position angle in the disk, which is not fully consistent with the field structures from the lens reconstruction, where ordered fields have only been observed in the north and east sides of the lensed galaxy. The compact size of SPT034652 favours a late-stage merger. It supports the possibility that galactic magnetic fields have already been established in the two merging galaxies in SPT034652, and the merger-triggered starburst makes the B-field more turbulent and leaves mainly the ordered magnetic fields in the outskirts, such as the north and east sides of the lensed galaxy in the source plane. Therefore, the galaxy merger provides the most natural explanation for the formation of ordered magnetic fields in SPT034652.
@@ 4.2.2 Outflow driven?
A strong outflow is present in SPT034652 based on the high-resolution dust continuum and H2O() observations (Jones et al. 2019a). It has a velocity up to 500 km s-1 and a substantial outflow rate, 100-900 M⊙ yr-1, likely driven by the galactic wind powered by the compact starbursting core. The galactic wind can distribute amplified magnetic fields to a larger volume (Brandenburg et al. 1993; Arámburo-García et al. 2021), thus can also explain the ordered field in the outskirts of SPT034652.lactic disks (Lopez-Rodriguez et al. 2021).
@@ 4.2.3 AGN driven?
It is also worth noting the possible contribution from the AGN in SPT034652. Powerful jets from AGN play an important role in magnetising the intracluster medium (ICM, Dubois et al. 2009; Xu et al. 2009). Meanwhile, the large-scale shocks induced by the jet and the AGN-driven outflow could also produce large-scale shear to order the turbulent fields. In the less extreme radio-quiet AGNs, the ubiquitous presence of compact radio jets (e.g. Girdhar et al. 2022; Rivera et al. 2023) also indicates their possible contribution to the development of magnetic fields in the ISM. In SPT034652, however, the presence of AGN is still elusive. The spectrum energy distribution (SED) analysis suggests a negligible AGN contribution to the total energy output, which is supported by the non-detection of X-ray and faint radio emission (Ma et al. 2016). In nearby AGN host galaxies, only the radio-loud AGN show highly polarised FIR fraction (5–11 per cent) in the core region; the radio-quiet AGN are mostly unpolarised (Lopez-Rodriguez et al. 2023). Therefore, even though there is a possible embedded AGN in SPT034652, our results support that it plays a negligible role in ordering the galactic magnetic fields. Future surveys targeting the AGN host galaxies are needed to test AGN’s importance in magnetising the ISM.
@@ 4.2.4 Caveats
In summary, a merger event provides the most likely explanation of the ordered magnetic fields in SPT034652, but we cannot fully rule out the possible contribution from outflows and a compact galactic disk. It is also possible that all these processes may work simultaneously to amplify and order the magnetic fields. Therefore, future high-resolution follow-ups will be the key to distinguishing different scenarios and alleviating the beam averaging.
@@ 4.3 Formation timescales of galactic magnetic fields
Theoretically, the amplification via small-scale dynamos is efficient, capable of amplifying the seed magnetic fields to G turbulent magnetic within a timescale of 100 Myr (Schleicher et al. 2010; Pakmor et al. 2014; Rieder & Teyssier 2016; Martin-Alvarez et al. 2022). However, the large-scale dynamos may need timescales of several Gyr to create kpc-scale ordered fields (e.g. Wang & Abel 2009; Kotarba et al. 2009; Arshakian et al. 2009; Rodrigues et al. 2019).
The starburst nature of SPT034652 allows efficient small-scale dynamos in the ISM. In local star-forming galaxies, the total magnetic field strength correlates with the SFR surface density, i.e. with to be around 0.3 (e.g. Heesen et al. 2014), indicating a close bond between the magnetic fields and star formation activities. In SPT034652, the SFR surface density is about . Such concentrated SFR normally indicates a highly turbulent ISM (Elmegreen & Scalo 2004; Bournaud et al. 2010), which may lead to the efficient amplification of turbulent magnetic fields. Meanwhile, the presence of turbulent magnetic fields has been used to connect the two SFR tracers: the total FIR luminosity and non-thermal radio luminosity (i.e. FIR-radio correlation) (Condon 1992; Yun et al. 2001; Bell 2003). The FIR-radio correlation has been observed up to by stacking analysis and, interestingly, it shows only mild variations with the redshift (e.g. Ivison et al. 2010; Magnelli et al. 2015; Delvecchio et al. 2021; Yoon 2024). This close correlation suggests that amplified magnetic fields may already be present in the star-forming galaxies up to , which supports the claim that small-scale dynamo efficiently amplifies the turbulent magnetic fields in the early star-forming galaxies. Fig. 4 shows the magnetic field strength as a function of redshift and total SFR. Due to the selection effects, the current two measurements of dust polarisation at , SPT0346-52 and 9io9, are both extreme starburst galaxies, which biased our results to the most extreme SFGs in the early Universe.
As we have mentioned before, the excess of Faraday rotation in the quasar absorber and fast radio burst support the presence of regular magnetic fields in galaxies up to (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Mannings et al. 2023). The detection of dust polarisation in 9io9 shows the ordered magnetic fields are present at z=2.6 (Geach et al. 2023). The redshift of SPT0346-52 is 5.6, thus the detection of the dust polarisation indicates the development of kpc-scale ordered galactic magnetic fields can be less than 1 Gyr, which puts a strong constraint on the ordering timescale. The current results do not allow us to directly test the dynamo. However, if the stretching and compressing during a merger is the main reason for the observed ordered magnetic fields (see the discussion in section 4.2), it indicates the two merging galaxies may have already developed ordered fields before the final merger, which indicates the ordering process is already ongoing during the first Gyr of the cosmic history. Therefore, our results suggest that the ordering process from large-scale dynamos could be efficient in early galaxies. Similar to the Faraday rotation, a more robust sample covering different galaxy types, redshifts, and environments is required to confirm this claim.
@@
where is the volume density of the ISM, and is the velocity dispersion of the gas. In SPT0346-52, the observed [C ii] line dispersion at the region with polarisation detection is about 140 km s-1 (see also Fig. 5). We assume a preferred gas density of 50 cm-3 from the photodissociation regions (PDR) modelling (Litke et al. 2022). We estimate the equipartition turbulent magnetic field strength to be G. We emphasise that this is only a theoretical upper limit for the turbulent magnetic fields, with the assumption that the turbulence dominates the line width. If the line width is dominated by regular rotation or the velocity difference of the possible two merging galaxies, the equipartition field strength can be several times lower, which will bring the field strength closer to the local starburst galaxies, either measured with dust polarisation (Lopez-Rodriguez et al. 2021; Lopez-Rodriguez 2023) or radio synchrotron observations (Lacki & Beck 2013). It should be noted that the magnetic field strength in the cold phase of the interstellar medium is different from that of the warm and diffuse medium measured at radio wavelengths (Martin-Alvarez et al. 2023). The field strength in the cold and dense medium could be larger than that from the warm and diffuse medium, as the field strength correlates with the gas volume density. Nonetheless, our calculation supports that highly amplified magnetic fields could be developed in very young star-forming galaxies.